# Reusing Historical Trajectories in Natural Policy Gradient via Importance Sampling: Convergence and Convergence Rate

### v1

@@ 6 Conclusion
In this paper, we study the convergence of a variant of natural policy gradient in reinforcement learning with reusing historical trajectories (RNPG). We provide a rigorous asymptotic convergence analysis of the RNPG algorithm by the ODE approach. Our results show that RNPG and its vanilla counterpart without reusing (VNPG) share the same limit ODE, while the bias resulting from the interdependence between iterations gradually diminishes, ultimately becoming insignificant in the asymptotic sense. We further demonstrate the benefit of reusing in RNPG and characterize the improved convergence rate of RNPG by the SDE approach. Through the numerical experiments on two classical benchmark problems, we verify our theoretical result and empirically study the choice of different reuse size in the RNPG algorithm.
@@ Acknowledgements
The authors are grateful for the support by Air Force Office of Scientific Research (AFOSR) under Grant FA9550-22-1-0244, National Science Foundation (NSF) under Grant DMS2053489, and Artificial Intelligence Institute for Advances in Optimization (AI4OPT) under Grant NSF-2112533. Yifan Lin and Yuhao Wang contributed equally to the manuscript.
@@ Appendix A Technical Proofs
A.1 Proof of Lemma 1
Proof.
First note that $\mathcal{F}_{n}$ provides all the information required to achieve $\theta_{n}$, we have $\theta_{n}\in\mathcal{F}_{n}$. Moreover, conditioned on $\mathcal{F}_{n}$, the expectation of the gradient estimator with historical trajectories take the following form
@@ A.1 Proof of Lemma 1
For $H(t)$, note that we add a small perturbation $\epsilon I_{d}$ to the FIM to ensure its positive definiteness to prevent the FIM from becoming singular. Hence, $\|\widehat{F}^{-1}(\theta_{n})\|\leq\epsilon^{-1}$ almost surely. Also note that the norm of $\widehat{\nabla\eta}(\theta_{n})$ is bounded w.p.1 under Assumption 1. Following the same argument in bounding the martingale difference sequence $\delta M_{n}$, we have $\sup_{n}\mathbb{E}\|\delta F_{n}\|<\infty$ and $\sup_{n}\mathbb{E}\|\delta F_{n}\|^{2}<\infty$, and thus $H(t)$ has zero asymptotic rate of change. ∎
A.2 Proof of Lemma 2
Proof.
For ease of notations, denote $\mathbb{E}_{n}[\cdot]=\mathbb{E}[\cdot|\boldsymbol{e}_{n}(\theta)=\boldsymbol{e}_{n},\theta]$. Note that $\forall\theta\in\Theta$, $\boldsymbol{e}_{i}(\theta)$ is independent of $\boldsymbol{e}_{n}(\theta)$ when $i\geq n+K-1$. So for those $i\geq n+K-1$, we have
@@ A.2 Proof of Lemma 2
for some constant $C_{3}>0$. Here we use the inequality $\|Mx\|\leq\|M\|\|x\|$, the fact that $\|M\|\leq\|M\|_{2}$ and $\|M\|_{2}$ is the largest eigenvalue for any positive definite matrix $M$ and vector $x$, under the condition (A.1.1), (A.1.2), (A.1.3) and (A.1.5) in Assumption 1. ∎
A.3 Proof of Lemma 3
Proof.
Recall that for $m\leq n+K-2$, we have
@@ A.3 Proof of Lemma 3
With Lemma 9, we have $\bar{F}(\theta)$ is Lipschitz continuous in $\theta$, thus $\bar{F}^{-1}(\theta)G(\xi;\theta)\frac{d^{\pi_{\theta}}(\xi)}{d^{\pi_{\theta^{\prime}}}(\xi)}$ is Lipschitz continuous in $\theta$, uniformly in $\xi$ and $n$. In addition, from Lemma 3.2 in [33], $\bar{F}^{-1}(\theta)\nabla\eta(\theta)$ is also Lipschitz continuous in $\theta$. Thus we prove that $v_{n}(\cdot,\boldsymbol{e})$ is Lipschitz continuous uniformly in $n$ and $\boldsymbol{e}$. One can then check that $b_{n}=O(\alpha_{n}^{2})$, which implies the asymptotic rate of change of $B(t)$ is zero under the condition (A.1.1) in Assumption 1. Following the same argument in Lemma 1, we have $I(t)$ has zero asymptotic rate of change. ∎
A.4 Proof of Corollary 1
Proof.
Relating the perturbed iteration to the original iteration, we have
@@ A.4 Proof of Corollary 1
We then have $\|Z(jT+t)-Z(jT)\|\leq\|B(jT+t)-B(jT)\|+\|I(jT+t)-I(jT)\|+\|v_{N(jT+t)}(\theta_{N(jT+t)})-\boldsymbol{e}_{N(jT+t)}\|+\|v_{N(jT)}(\theta_{N(jT)})-\boldsymbol{e}_{N(jT)}\|$. By Lemma 2 and Lemma 3, $Z(t)$ has zero asymptotic rate of change w.p.1. ∎
A.5 Proof of Corollary 2
Proof.
For any $s>0$, let $\boldsymbol{\xi}_{s}:=(\xi_{s}^{1},\cdots,\xi_{s}^{B})$, $\boldsymbol{\pi}_{s}:=(\pi_{\theta}(\xi_{s}^{1}),\cdots,\pi_{\theta}(\xi_{s}^{B}))$, effective memory $\boldsymbol{e}_{s}:=(\boldsymbol{\xi}_{s-K+1}$, $\boldsymbol{\pi}_{s-K+1},\cdots,\boldsymbol{\xi}_{s-1},\boldsymbol{\pi}_{s-1})$, and non-decreasing filtration $\mathcal{F}_{n}:=\sigma\{(\theta_{s},\boldsymbol{e}_{s}),s\leq n\}$. The rest of the proof follows by replacing the occupancy measure $d^{\pi_{\theta}}$ by the policy $\pi_{\theta}$. Note that there is a slight modification of (A.1.5) in Assumption 1: there exists a constant $0<\epsilon_{\pi}\leq 1$ such that the policy $\pi_{\theta}(s,a)\geq\epsilon_{\pi},\forall(s,a)\in\mathcal{S}\times\mathcal{A},\forall\theta\in\Theta$. ∎
A.6 Proof of Lemma 4
Proof.
Denote $\bar{F}^{-1}_{n}=\mathbb{E}_{n}[\widehat{F}^{-1}_{n}]$. For $i\geq n+K$, we have
@@ A.6 Proof of Lemma 4
The second equality holds because of the tower property and that conditioned on $\theta_{i}$, $\bar{F}^{-1}_{i}$ is a constant. Next, since $\theta_{i}\rightarrow\bar{\theta}$ w.p.1, $\bar{F}^{-1}(\theta)$ is continuous in $\theta$ and $\Theta$ is compact, by bounded convergence theorem we have $\mathbb{E}[L_{i}L_{i}^{T}]\rightarrow\frac{1}{KB^{2}}\Sigma_{1}(\bar{\theta})$ w.p.1. Since the first $K$ terms do not affect the average value as $m$ goes to infinity, we complete the proof. ∎
A.7 Proof of Lemma 5
Proof.
For any $i\geq n+K$, we have
@@ A.7 Proof of Lemma 5
Furthermore, since the first $K$ terms do not affect the average value when $m$ goes to infinity, we complete the proof. ∎
A.8 Proof of Lemma 6
Proof.
First, note that $\mathbb{E}_{n}\left[L_{i}R_{i}^{T}\right]=\mathbb{E}_{n}\left[\left(\widehat{F}^{-1}_{i}-\mathbb{E}_{i}[\widehat{F}^{-1}_{i}]\right)\widehat{\nabla\eta}_{i}\left(\mathbb{E}_{i}[\widehat{F}_{i}^{-1}]\left(\widehat{\nabla\eta}_{i}-\mathbb{E}_{i}[\widehat{\nabla\eta}_{i}]\right)\right)^{T}\right]$. Since conditioned on $\theta_{i},\boldsymbol{e}_{i}$, $\widehat{F}_{i}^{-1}$ is independent of $\widehat{\nabla\eta}_{i}$, we can obtain
@@ A.8 Proof of Lemma 6
Hence, the Lemma holds. ∎
A.9 Proof of Lemma 7
Proof.
For $i\geq n+K$, we have
@@ A.9 Proof of Lemma 7
Since the first $K$ terms does not affect the limit, we obtain the desired result. ∎
A.10 Proof of Lemma 8
Proof.
Note that when $i\geq n+K$, $\boldsymbol{e}_{i}(\theta)$ is independent of $\boldsymbol{e}_{n}(\theta)$. So, we have
@@ A.10 Proof of Lemma 8
Finally, since the first $K$ ($i<n+K$) terms do not affect the limit, we complete the proof. ∎
A.11 Proof of Theorem 3
Proof.
Let $\mathcal{L}_{n}=\nabla\eta(\theta^{*})-\frac{1}{B}\sum_{i=1}^{B}G(\xi_{n}^{i},\theta^{*},\theta_{n})$. Let $\mathcal{F}^{\prime}_{n}=\sigma\{(\theta_{m},\boldsymbol{\xi}_{m}),m\leq n\}$. Then we have $\frac{1}{B}\sum_{i=1}^{B}G(\xi_{n}^{i},\theta^{*},\theta_{n})$ is a conditional unbiased estimator of $\nabla\eta(\theta^{*})$, given the information of past $n-1$ iterations. Then $(\mathcal{L}_{n},\mathcal{F}^{\prime}_{n})_{n\geq 0}$ is a martingale difference sequence, and $(\sum_{m=n-K+1}^{s}\mathcal{L}_{s},\mathcal{F}^{\prime}_{s})_{s=n-K+1}^{n}$ is a martingale sequence. We can then write the gradient estimation error in terms of $\mathcal{L}$:
$\nabla\eta(\theta^{*})-\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}G(\xi_{m}^{i},\theta^{*},\theta_{m})=\frac{1}{K}\sum_{m=nn-K+1}^{n}\mathcal{L}_{m}.$
To bound the gradient estimation error, we first introduce the Freedman’s inequality for matrix martingale.
Lemma 10 (Corollary 1.3 in [35]).
Let $(\mathcal{L}_{i},\mathcal{F}_{i})_{i=1,\cdots,n}$ be a martingale difference sequence. Suppose $\|\mathcal{L}_{i}\|_{2}\leq C$. Then for all $\epsilon,\nu>0$, we have
$P\left(\|\sum_{i=1}^{k}\mathcal{L}_{i}\|_{2}\geq\epsilon\text{ and }\sum_{i=1}^{k}\mathbb{E}[\|\mathcal{L}_{i}\|_{2}^{2}|\mathcal{F}_{i-1}\leq\nu^{2}\text{ for some }k\leq n]\right)\leq(d+1)\exp\left(-\frac{\epsilon^{2}}{2(\nu^{2}+\epsilon C)}\right)$
By Lemma 10, we have
$P\left(\|\sum_{m=n-K+1}^{n}\mathcal{L}_{m}\|_{2}\geq\epsilon\text{ and }\sum_{m=n-K+1}^{n}\mathbb{E}[\|\mathcal{L}_{m}\|_{2}^{2}|\mathcal{F}_{m-1}]\leq\nu^{2}\right)\leq(d+1)\exp\left(-\frac{\epsilon^{2}}{2(\nu^{2}+\epsilon C)}\right).$ (13)
Note that $\mathbb{E}[\|\mathcal{L}_{m}\|_{2}^{2}|\mathcal{F}^{\prime}_{m-1}]=\frac{1}{B}\operatorname{trace}(\operatorname{Var}_{\xi\sim d^{\pi_{\theta_{m}}}}[G(\xi,\theta^{*},\theta_{m})|\mathcal{F}^{\prime}_{m-1}])$, $\mathbb{E}[\widehat{\operatorname{Var}}_{m}|\mathcal{F}^{\prime}_{m-1}]=\operatorname{Var}_{\xi\sim d^{\pi_{\theta_{m}}}}[G(\xi,\theta^{*},\theta_{m})|\mathcal{F}^{\prime}_{m-1}]$, and $0\leq\operatorname{trace}(\widehat{\operatorname{Var}}_{m})\leq\frac{B}{B-1}C^{2}$. Applying McDiarmid’s inequality (see [36]), we have
$P\left(\sum_{m=n-K+1}^{n}(\frac{1}{B}\operatorname{trace}(\widehat{\operatorname{Var}}_{m})-\mathbb{E}[\|\mathcal{L}_{m}\|_{2}^{2}|\mathcal{F}^{\prime}_{m-1}])\geq\epsilon\right)\leq\exp\left(-\frac{B(B-1)\epsilon^{2}}{KC^{2}}\right).$
For any $\delta>0$, take $\epsilon=\sqrt{\frac{KC^{2}}{B(B-1)}\log\frac{\pi^{2}n^{2}}{3\delta}}$, $\beta^{\prime}_{n}=\epsilon+\frac{1}{B}\sum_{m=n-K+1}^{n}\operatorname{trace}(\widehat{\operatorname{Var}}_{m})$, then
$P\left(\sum_{m=n-K+1}^{n}\mathbb{E}[\|\mathcal{L}_{m}\|^{2}_{2}|\mathcal{F}^{\prime}_{m-1}]\geq\beta^{\prime}_{n}\right)\leq\frac{3\delta}{\pi^{2}n^{2}}.$
In order the right hand side of (13) to be smaller than or equal to $\frac{3\delta}{\pi^{2}n^{2}}$ given $\nu^{2}=\beta^{\prime}_{n}$, i.e., $(d+1)\exp\left(-\frac{\epsilon^{\prime 2}}{2(\beta^{\prime}_{n}+\epsilon^{\prime}C)}\right)\leq\frac{3\delta}{\pi^{2}n^{2}}$, we require
$\epsilon^{\prime}\geq C\log\left(\frac{\pi^{2}n^{2}(d+1)}{3\delta}\right)+\sqrt{C^{2}\left(\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}\right)^{2}+2\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}\beta^{\prime}_{n}}.$
Using the inequality $\sqrt{a^{2}+b^{2}}\leq a+b$ for positive $a$ and $b$, we can take $\epsilon^{\prime}=2C\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}+\sqrt{2\beta^{\prime}_{n}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}}$. Then we have
$P\left(\left\|\sum_{m=n-K+1}^{n}\mathcal{L}_{m}\right\|_{2}\geq\epsilon^{\prime}\right)$
$=$ $P\left(\left\|\sum_{m=n-K+1}^{n}\mathcal{L}_{m}\right\|_{2}\geq\epsilon^{\prime}\text{ and }\sum_{m=n-K+1}^{n}\mathbb{E}[\left\|\mathcal{L}_{m}\right\|^{2}_{2}|\mathcal{F}^{\prime}_{m-1}]\leq\beta^{\prime}_{n}\right)$
$+P\left(\left\|\sum_{m=n-K+1}^{n}\mathcal{L}_{m}\right\|_{2}\geq\epsilon^{\prime}\text{ and }\sum_{m=n-K+1}^{n}\mathbb{E}[\|\mathcal{L}_{m}\|^{2}_{2}|\mathcal{F}^{\prime}_{m-1}]\geq\beta^{\prime}_{n}\right)$
$=$ $P\left(\left\|\sum_{m=n-K+1}^{n}\mathcal{L}_{m}\right\|_{2}\geq\epsilon^{\prime}\text{ and }\sum_{m=n-K+1}^{n}\mathbb{E}[\|\mathcal{L}_{m}\|^{2}_{2}|\mathcal{F}^{\prime}_{m-1}]\leq\beta^{\prime}_{n}\right)$
$+P\left(\sum_{m=n-K+1}^{n}\mathbb{E}[\|\mathcal{L}_{m}\|^{2}_{2}|\mathcal{F}^{\prime}_{m-1}]\geq\beta^{\prime}_{n}\right)$
$\leq$ $(d+1)\exp\left(-\frac{\epsilon^{\prime 2}}{2(\beta^{\prime}_{n}+\epsilon^{\prime}C)}\right)+\frac{3\delta}{\pi^{2}n^{2}}$
$\leq$ $\frac{3\delta}{\pi^{2}n^{2}}+\frac{3\delta}{\pi^{2}n^{2}}=\frac{6\delta}{\pi^{2}n^{2}}.$
Therefore, with probability at least $1-\frac{6\delta}{\pi^{2}n^{2}}$, we have $\|\sum_{m=n-K+1}^{n}\mathcal{L}_{m}\|_{2}\leq\left(2C\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}+\sqrt{2\beta^{\prime}_{n}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}}\right)$. Let $\beta_{n}=\frac{\beta^{\prime}_{n}}{K^{2}}$, i.e., $\beta_{n}=\sqrt{\frac{C^{2}}{K^{3}B(B-1)}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}}+\frac{1}{K^{2}B}\sum_{m=n-K+1}^{n}\operatorname{trace}(\widehat{\operatorname{Var}}_{m})$, we have with probability at least $1-\frac{6\delta}{\pi^{2}n^{2}}$,
$\left\|\nabla\eta(\theta^{*})-\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}G(\xi_{m}^{i},\theta^{*},\theta_{m})\right\|_{2}=\left\|\frac{1}{K}\sum_{m=n-K+1}^{n}\mathcal{L}_{m}\right\|_{2}$
$\leq$ $\sqrt{2\beta^{\prime}_{n}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}}+\frac{2C}{K}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}.$ (14)
Denote $\mathcal{E}_{n}$ as the event that inequality (A.11) holds for $n$.
$P(\cap_{n=1}^{\infty}\mathcal{E}_{n})$ $=1-P(\cup_{n=1}^{\infty}\mathcal{E}_{n}^{c})$
$\geq 1-\sum_{n=1}^{\infty}P(\mathcal{E}_{n}^{c})$
$\geq 1-\sum_{n=1}^{\infty}\frac{6\delta}{\pi^{2}n^{2}}$
$=1-\delta,$
where the first inequality uses the union bound. Therefore, we have with probability at least $1-\delta$,
$\left\|\nabla\eta(\theta^{*})-\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}G(\xi_{m}^{i},\theta^{*},\theta_{m})\right\|_{2}$
$\leq\sqrt{2\beta_{n}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}}+\frac{2C}{K}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}.$

### v2

@@ 6 Conclusion
In this paper, we study the convergence of a variant of natural policy gradient in reinforcement learning with reusing historical trajectories (RNPG). We provide a rigorous asymptotic convergence analysis of the RNPG algorithm by the ODE approach. Our results show that RNPG and its vanilla counterpart without reusing (VNPG) share the same limit ODE, while the bias resulting from the interdependence between iterations gradually diminishes, ultimately becoming insignificant in the asymptotic sense. We further demonstrate the benefit of reusing in RNPG and characterize the improved convergence rate of RNPG by the SDE approach. Through numerical experiments on two classical benchmark problems, we demonstrate the performance of RNPG and reusing samples in proximal policy optimization (PPO), and empirically study the choice of different reuse sizes. We also empirically verify the theoretical asymptotic normality result on a linear quadratic control problem.
@@ Acknowledgements
The authors are grateful for the support by Air Force Office of Scientific Research (AFOSR) under Grant FA9550-22-1-0244, National Science Foundation (NSF) under Grant ECCS-2419562, and Artificial Intelligence Institute for Advances in Optimization (AI4OPT) under Grant NSF-2112533. The first author (Yifan Lin) and second author (Yuhao Wang) contributed equally to the manuscript.
Code and Data
The code used for the experiments in Section 5 is available in the online supplementary materials.
@@ Appendix A Technical Proofs
A.1 Sufficient condition for (A.2.3)
Lemma 9
Suppose $\eta(\theta)$ is locally strongly concave at $\bar{\theta}$. Then (A.2.3) holds true.
Proof. Since $\eta(\theta)$ is locally strongly concave at $\bar{\theta}$, there exists $\mu>0$, such that $\nabla^{2}\eta(\theta)\preccurlyeq-\mu I_{d}$ for all $\theta$ in a small neighborhood of $\bar{\theta}$. Then, we can write
$\nabla\eta(\theta)=\nabla\eta(\bar{\theta})+\nabla^{2}\eta(\bar{\theta})(\theta-\bar{\theta})+o(\|\theta-\bar{\theta}\|)=\nabla^{2}\eta(\bar{\theta})(\theta-\bar{\theta})++o(\|\theta-\bar{\theta}\|).$
Then
$\bar{F}^{-1}(\theta)\nabla\eta(\theta)$
$=$ $\bar{F}^{-1}(\theta)\nabla^{2}\eta(\bar{\theta})(\theta-\bar{\theta})+o(\|\theta-\bar{\theta}\|)$
$=$ $\bar{F}^{-1}(\bar{\theta})\nabla^{2}\eta(\bar{\theta})(\theta-\bar{\theta})+(\bar{F}^{-1}(\theta)-\bar{F}^{-1}(\bar{\theta}))\nabla^{2}\eta(\bar{\theta})(\theta-\bar{\theta})+o(\|\theta-\bar{\theta}\|)$
With regularity the assumption on $\pi(\theta)$ (Lipschitz continuity, A.1.3), we have (by Proof of Lemma 2) $(\bar{F}^{-1}(\theta)-\bar{F}^{-1}(\bar{\theta}))=O(\|\theta-\bar{\theta}\|)$. This leads to
$\bar{F}^{-1}(\theta)\nabla\eta(\theta)=\bar{F}^{-1}(\bar{\theta})\nabla^{2}\eta(\bar{\theta})(\theta-\bar{\theta})+o(\|\theta-\bar{\theta}\|).$
Since
$\bar{F}^{-1}(\bar{\theta})=\mathbb{E}\left[\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{i},\bar{\theta})\right)^{-1}\right],$
$\bar{F}^{-1}(\bar{\theta})$ is positive definite. Furthermore, we know $\nabla^{2}\eta(\bar{\theta})\preccurlyeq-\mu I_{d}$ negative definite. To show $\bar{F}^{-1}(\theta)\nabla\eta(\theta)$ is Hurwitz, it is sufficient to show $\bar{F}^{-1}(\theta)\nabla\eta(\theta)$ has $d$ negative eigenvalues. Notably, $\bar{F}^{-1}(\theta)\nabla\eta(\theta)$ is not necessarily negative definite since it may not be symmetric. In the following we show for any positive definite matrix $A$ and negative definite matrix $B$, $A,B\in\mathbb{R}^{d\times d}$, the multiplication $AB$ has $d$ negative eigenvalues.
Since $A$ is positive definite, there exists positive definite matrix $P$, such that $P^{2}=A$. Then,
$AB=P^{2}B=P(PBP)P^{-1}.$
That is, $M:=PBP$ is a similar matrix for $AB$, and $M$ has the same eigenvalues as $AB$. Note $M^{\top}=P^{\top}B^{\top}P^{\top}=PBP=M$. $M$ is symmetric. Further notice $\forall x\in\mathbb{R}^{d}$ and $x\neq\mathbf{0}$,
$x^{\top}Mx=x^{\top}PBPx=x^{\top}P^{\top}BPx=(Px)^{\top}B(Px),$
since $P$ is positive definite, $Px\neq\mathbf{0}$. Furthermore, since $B$ is negative definite, $(Px)^{\top}B(Px)=x^{\top}Mx<0$. This implies $M$ is negative definite, and hence, $M$ has $d$ negative eigenvalues. This implies $AB$ also has $d$ negative eigenvalues, which means that $AB$ is a Hurwitz matrix. $\blacksquare$
A.2 Proof of Lemma 1
Proof. First note that $\mathcal{F}_{n}$ provides all the information required to achieve $\theta_{n}$, we have $\theta_{n}\in\mathcal{F}_{n}$. Moreover, conditioned on $\mathcal{F}_{n}$, the expectation of the gradient estimator with historical trajectories take the following form
@@ A.2 Proof of Lemma 1
For $H(t)$, note that we add a small perturbation $\epsilon I_{d}$ to the FIM to ensure its positive definiteness to prevent the FIM from becoming singular. Hence, $\|\widehat{F}^{-1}(\theta_{n})\|\leq\epsilon^{-1}$ almost surely. Also note that the norm of $\widehat{\nabla\eta}(\theta_{n})$ is bounded w.p.1 under Assumption 1. Following the same argument in bounding the martingale difference sequence $\delta M_{n}$, we have $\sup_{n}\mathbb{E}\|\delta F_{n}\|<\infty$ and $\sup_{n}\mathbb{E}\|\delta F_{n}\|^{2}<\infty$, and thus $H(t)$ has zero asymptotic rate of change. $\blacksquare$
A.3 Proof of Lemma 2
Proof. For ease of notations, denote $\mathbb{E}_{n}[\cdot]=\mathbb{E}[\cdot|\boldsymbol{e}_{n}(\theta)=\boldsymbol{e}_{n},\theta]$. Note that $\forall\theta\in\Theta$, $\boldsymbol{e}_{i}(\theta)$ is independent of $\boldsymbol{e}_{n}(\theta)$ when $i\geq n+K-1$. So for those $i\geq n+K-1$, we have
@@ A.3 Proof of Lemma 2
for some constant $C_{3}>0$. Here we use the inequality $\|Mx\|\leq\|M\|\|x\|$, the fact that $\|M\|\leq\|M\|_{2}$ and $\|M\|_{2}$ is the largest eigenvalue for any positive definite matrix $M$ and vector $x$, under the condition (A.1.1), (A.1.2), (A.1.3) and (A.1.5) in Assumption 1. $\blacksquare$
A.4 Proof of Lemma 3
Proof. Recall that for $m\leq n+K-2$, we have
@@ A.4 Proof of Lemma 3
$\blacksquare$
With Lemma 10, we have $\bar{F}(\theta)$ is Lipschitz continuous in $\theta$, thus $\bar{F}^{-1}(\theta)G(\xi;\theta)\frac{d^{\pi_{\theta}}(\xi)}{d^{\pi_{\theta^{\prime}}}(\xi)}$ is Lipschitz continuous in $\theta$, uniformly in $\xi$ and $n$. In addition, from Lemma 3.2 in [34], $\bar{F}^{-1}(\theta)\nabla\eta(\theta)$ is also Lipschitz continuous in $\theta$. Thus we prove that $v_{n}(\cdot,\boldsymbol{e})$ is Lipschitz continuous uniformly in $n$ and $\boldsymbol{e}$. One can then check that $b_{n}=O(\alpha_{n}^{2})$, which implies the asymptotic rate of change of $B(t)$ is zero under the condition (A.1.1) in Assumption 1. Following the same argument in Lemma 1, we have $I(t)$ has zero asymptotic rate of change. $\blacksquare$
A.5 Proof of Corollary 1
Proof. Relating the perturbed iteration to the original iteration, we have
@@ A.5 Proof of Corollary 1
We then have $\|Z(jT+t)-Z(jT)\|\leq\|B(jT+t)-B(jT)\|+\|I(jT+t)-I(jT)\|+\|v_{N(jT+t)}(\theta_{N(jT+t)})-\boldsymbol{e}_{N(jT+t)}\|+\|v_{N(jT)}(\theta_{N(jT)})-\boldsymbol{e}_{N(jT)}\|$. By Lemma 2 and Lemma 3, $Z(t)$ has zero asymptotic rate of change w.p.1. $\blacksquare$
A.6 Proof of Corollary 2
Proof. For any $s>0$, let $\boldsymbol{\xi}_{s}:=(\xi_{s}^{1},\cdots,\xi_{s}^{B})$, $\boldsymbol{\pi}_{s}:=(\pi_{\theta}(\xi_{s}^{1}),\cdots,\pi_{\theta}(\xi_{s}^{B}))$, effective memory $\boldsymbol{e}_{s}:=(\boldsymbol{\xi}_{s-K+1}$, $\boldsymbol{\pi}_{s-K+1},\cdots,\boldsymbol{\xi}_{s-1},\boldsymbol{\pi}_{s-1})$, and non-decreasing filtration $\mathcal{F}_{n}:=\sigma\{(\theta_{s},\boldsymbol{e}_{s}),s\leq n\}$. The rest of the proof follows by replacing the occupancy measure $d^{\pi_{\theta}}$ by the policy $\pi_{\theta}$. Note that there is a slight modification of (A.1.5) in Assumption 1: there exists a constant $0<\epsilon_{\pi}\leq 1$ such that the policy $\pi_{\theta}(s,a)\geq\epsilon_{\pi},\forall(s,a)\in\mathcal{S}\times\mathcal{A},\forall\theta\in\Theta$. $\blacksquare$
A.7 Proof of Lemma 4
Proof. Denote $\bar{F}^{-1}_{n}=\mathbb{E}_{n}[\widehat{F}^{-1}_{n}]$. For $i\geq n+K$, we have
@@ A.7 Proof of Lemma 4
The second equality holds because of the tower property and that conditioned on $\theta_{i}$, $\bar{F}^{-1}_{i}$ is a constant. Next, since $\theta_{i}\rightarrow\bar{\theta}$ w.p.1, $\bar{F}^{-1}(\theta)$ is continuous in $\theta$ and $\Theta$ is compact, by bounded convergence theorem we have $\mathbb{E}[L_{i}L_{i}^{T}]\rightarrow\frac{1}{KB^{2}}\Sigma_{1}(\bar{\theta})$ w.p.1. Since the first $K$ terms do not affect the average value as $m$ goes to infinity, we complete the proof. $\blacksquare$
A.8 Proof of Lemma 5
Proof. For any $i\geq n+K$, we have
@@ A.8 Proof of Lemma 5
Furthermore, since the first $K$ terms do not affect the average value when $m$ goes to infinity, we complete the proof. $\blacksquare$
A.9 Proof of Lemma 6
Proof. First, note that $\mathbb{E}_{n}\left[L_{i}R_{i}^{T}\right]=\mathbb{E}_{n}\left[\left(\widehat{F}^{-1}_{i}-\mathbb{E}_{i}[\widehat{F}^{-1}_{i}]\right)\widehat{\nabla\eta}_{i}\left(\mathbb{E}_{i}[\widehat{F}_{i}^{-1}]\left(\widehat{\nabla\eta}_{i}-\mathbb{E}_{i}[\widehat{\nabla\eta}_{i}]\right)\right)^{T}\right]$. Since conditioned on $\theta_{i},\boldsymbol{e}_{i}$, $\widehat{F}_{i}^{-1}$ is independent of $\widehat{\nabla\eta}_{i}$, we can obtain
@@ A.9 Proof of Lemma 6
Hence, the Lemma holds. $\blacksquare$
A.10 Proof of Lemma 7
Proof. For $i\geq n+K$, we have
@@ A.10 Proof of Lemma 7
Since the first $K$ terms does not affect the limit, we obtain the desired result. $\blacksquare$
A.11 Proof of Lemma 8
Proof. Note that when $i\geq n+K$, $\boldsymbol{e}_{i}(\theta)$ is independent of $\boldsymbol{e}_{n}(\theta)$. So, we have
@@ A.11 Proof of Lemma 8
Finally, since the first $K$ ($i<n+K$) terms do not affect the limit, we complete the proof. $\blacksquare$